# Threshold Graphs are Globally Synchronizing

### v1

@@ 2. Preliminaries of threshold graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines. They were in fact discovered independently and reported in different journals by researchers working in distinct fields. The term “threshold graph” was introduced by Chvátal and Hammer in their study of set-packing problems. Around the same time, the same class of graphs was independently discovered by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs. Since then, threshold graphs have appeared in various contexts, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences.
Many equivalent conditions have been found for threshold graphs including constructive forms and forbidden configurations. See for example [MP95] for nine different characterizations.
@@
A graph is threshold if it contains no induced $P_{4}$, $C_{4}$ and $2K_{2}$.
One of the many characterizations of threshold graphs is constructive. The construction is in terms of two basic operations; adding a disjoint vertex (adjacent to no existing vertex), or a dominating vertex (adjacent to all existing vertices).
The construction process can be encoded into a binary sequence, called threshold sequence, that is $(b_{1},\dots,b_{n})\in\{0,1\}^{n}$ such that starting from a single vertex, for each $b_{i}=0$ we add an isolated vertex, and for each $b_{i}=1$ we add a dominating vertex; see Figure 1.
@@
Figure 2. Forbidden induced subgraphs for threshold graphs: graphs are $\{P_{4},C_{4},2K_{2}\}$-free. Only edges within each induced pattern are drawn.
@@
Interestingly, threshold graphs have also been studied in synchronization contexts beyond oscillator models. In particular, [PS95] surveyed their use in the design of synchronization protocols for cooperating processes, focusing on mutual exclusion via semaphore-based control mechanisms and the so-called PV-chunk semaphores. In some papers, these are also referred to as Pvc-definable graphs.
3. Preliminaries on Kuramoto synchronization
In this section, we aim to establish the correspondence between dynamical and optimization concepts in the homogeneous Kuramoto model. Since the homogeneous Kuramoto model (2) can be formulated as a gradient dynamical system with the energy function (3) and the real analycity of (3), the dynamical notions of equilibrium points and stable equilibria naturally correspond to the optimization notions of first-order and second-order stationary points of the energy function. Before showing this correspondence, we first provide the formal definitions of (stable) equilibria in dynamical systems and of first- and second-order stationary points in optimization. It is worth noting that such a one-to-one correspondence between the two sets of concepts is rather uncommon. This special structure is precisely what allows the analysis of global synchronization in the Kuramoto model to be directly connected to the geometry of the nonconvex optimization landscape.
In general, the evolution of a dynamical system can often be viewed as a descent process of a potential, known as a Lyapunov function, which monotonically decreases along the system trajectories. However, Lyapunov functions are not unique in general, let alone the existence of a clear correspondence between the two classes of concepts discussed above. Moreover, even if the system is of gradient type so that its energy function serves as a natural candidate for a Lyapunov function, such a correspondence is still not guaranteed in general. The remarkable property of the homogeneous Kuramoto model considered here is that its energy function is real analytic, and it is precisely this property that ensures the correspondence between dynamical and optimization notions. For further discussion and proofs, we refer the reader to [AK06].
@@ 3.1. Equilibrium and local stability
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@ 3.2. Stationary point and local optimality
3.3. Correspondence between stability and geometry
@@
In this section, we briefly outline the relationship between equilibrium stability in the dynamical system (2) and the geometric structure of the energy function (3). Understanding this correspondence provides valuable intuition. In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order critical point.
@@
At the first-order level, the two notions coincide: a configuration $\boldsymbol{\theta}$ is a critical point of the energy function $E(\boldsymbol{\theta})$ if and only if it is an equilibrium of the dynamical system (2), since both are characterized by
$\sum_{j=1}^{n}A_{ij}\sin(\theta_{i}-\theta_{j})=0\text{ for all }i\in[n]$
@@
One can further relate the stability of an equilibrium to second-order conditions of the energy landscape. It is precisely because of the real analycity of the energy function (3), the local asymptotic stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (3). This is a remarkable property since it does not hold in general for dynamical systems. In general, the dynamics of a system may admit more complex behaviors near equilibria such as limit cycles, or chaotic dynamics, etc. The relation mentioned here is summarized in Figure 3. One can find the proof in the celebrated work [AK06].
SOSP of 3 Local minima of 3 = stable equilibria of 2
Figure 3. Relation between different concepts. Here equivalence between stable equilibrium and local minimum is established in [AK06].
3.4. Synchronous and aligned states
In this section, we prove that all aligned states are equilibria. Among them, however, the non-synchronous ones are not second-order stationary points (SOSPs). Before that, we first define the notions of synchronous states and aligned states.
Definition 3.6.
A state $(\theta_{1},\cdots,\theta_{n})$ of $G=(V,E)$ is called a synchronous state if for all $i,j\in V$,
$\theta_{i}\equiv\theta_{j}\pmod{2\pi}$.
More generally, it is called an aligned state if for all $i,j\in V$,
$\theta_{i}\equiv\theta_{j}\pmod{\pi}$.
Clearly, every synchronous state is also an aligned state. The stability of all aligned states is:
Theorem 3.7.
Any aligned state is an equilibrium of the Kuramoto model (2) on the graph $G$ or equivalently stationary point of $E(\boldsymbol{\theta})$. Among them, synchronous states are second-order critical point of (3) [^2] In fact, they are asymptotically stable equilibria up to rotational invariance, while non-synchronous aligned states are not second-order critical point of (3).
Proof.
Firstly, it is straightforward to see that aligned states satisfy the condition (4), therefore they are equilibria of $G$, or equivalently stationary point of $E(\boldsymbol{\theta}$.
Secondly, we turn to the second-order criticality of synchronous states. The Hessian matrix at synchronous states is given by
$\boldsymbol{\mathbf{H}}_{ij}=\begin{cases}\sum_{j=1}^{n}\boldsymbol{A}_{ij},&\text{if }i=j,\\[8.0pt] -\boldsymbol{A}_{ij},&\text{if }i\neq j,\end{cases}$ (6)
[…]
Thirdly, regarding the second-order criticality of non-synchronous aligned states. Without loss of generality [^3] Any aligned state can be rotated uniformly on all angles $\theta_{1},\cdots,\theta_{n}$, such that $\theta_{i}=0$ or $\pi$, for any $i\in[n]$., assume $\theta_{i}=0$ or $\pi$, for any $i\in[n]$. Let $\boldsymbol{x}=(x_{1},\cdots,x_{n})$ be an $n$-dimensional vector that indicates which group (either at angle 0 or $\pi$) each node belongs to
[…]
Then the Hessian matrix at any antipodal state is not negative semidefiniteness. Indeed, this can be inferred by the quadratic form
$\boldsymbol{x}^{T}\boldsymbol{\mathbf{H}}\boldsymbol{x}=\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=-\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}<0$
since $\boldsymbol{A}_{ij}\geq 0$, and there must exist at least one pair of nodes $i,j$ such that $\theta_{x}\equiv\theta_{y}\pmod{\pi}$, $k\in\mathbb{Z}$, given the connectness of $G$. Therefore, any non-synchronous aligned state is not a SOSP. ∎
Remark 3.8.
In fact, synchronous states are asymptotically stable equilibria up to a global rotation. Here, asymptotic stability is a stronger notion than mere stability: it requires that small perturbations lead to trajectories returning to the equilibrium. On the other hand, non-synchronous aligned states are not SOSPs and therefore cannot be local minima of 3 or say not stable equilibrium of 2.
The instability of non-synchronous states can be intuitively understood as follows: a set of antipodally positioned oscillators must evolve in a way that decreases the system’s energy, as there always exists an interaction force between these antipodal pairs that drives them closer together.
$A_{12}1122$
Figure 4. Synchronous (left) and non-synchronous aligned state (right). Arrows indicate the nodes moving closer along the circle, implying this state is not stable.
4. Combinatorial planar geometry of equilibrium
As we will introduce later, the key idea of our proof is to assume that a configuration $\boldsymbol{\theta}$ is a second-order critical point and show that this assumption forces $\boldsymbol{\theta}$ to be synchronized, meaning that all node vectors coincide. This requires us to understand the geometric configuration of second-order critical points on the unit circle. In the previous section, we presented the algebraic conditions characterizing first- and second-order critical points.
In this section, we employ a complex-number formulation to transform these algebraic conditions into geometric ones, i.e., combinatorial relations between the vectors associated with the nodes. This seemingly simple transformation significantly simplifies the original landscape geometry problem, reducing it to a combinatorial-geometric problem involving vectors on the two-dimensional plane. With this, one just needs to understand whether, under the geometric form of the first- and second-order conditions, it can be deduced that all vectors must be identical.
4.1. Vector-labelled graph
@@
Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$, we associate to each node $i\in V$ its phasor
@@
This yields a vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ which encodes the phase configuration $\boldsymbol{\theta}$ geometrically.
$v_{i}$$\theta_{i}$
Figure 5. Representation of $v_{i}$ with angle $\theta_{i}$.
Although straightforward to construct, this object captures both the equilibrium condition and its stability in the Kuramoto model (2) on the graph $G$.
More precisely, to determine whether a given $\boldsymbol{\theta}$ is an equilibrium and whether it is stable (or more general, second-order critical point), the combinatorial geometry of the vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ provides a geometric view to algebraic one (solving nonlinear equations and analyzing Hessian at stationary point), as detailed in Section 4.2.
@@ 4.2. Alignment and equilibrium
Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms. For each node $i\in V$, we aggregate the phasors of its neighbors into a single vector:
$\sum_{j\in N(i)}\boldsymbol{v}_{j}.$
This vector represents the aggregated state of node $i$’s neighbors. If it is not aligned with $\boldsymbol{v}_{i}$, then intuitively, due to the intrinsic synchronization mechanism in the Kuramoto model, the nodes will tend to move toward alignment instead of staying there at rest, therefore not in equilibrium. The equivalence between the alignment $\sum_{j\in N(i)}\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}$ for every node $i\in V$ and the equilibrium condition can be rigorously established by representing the phasors as complex numbers.
@@
The expression describes a geometric condition of equilibrium: At equilibrium, each phasor aligns with the direction of the sum of its neighbors’ vectors and is scaled to unit length. Although introduced as a normalization factor, the quantity $\mu_{i}$ carries physical meanings. If the neighbors are highly coherent, the sum vector is large, requiring a larger $\mu_{i}$ to normalize $\boldsymbol{v}_{i}$ to unit length; if the directions are dispersed, $\mu_{i}$ is small. In this sense, $\mu_{i}$ measures how strongly the phasors of node $i$’s neighbors align directionally. Indeed, defining the local order parameter
$R_{i}:=\frac{1}{\deg(i)}\left|\sum_{j\in N(i)}e^{i\theta_{j}}\right|,$
we have the relation $\mu_{i}=\deg(i)\cdot R_{i}$, showing that $\mu_{i}$ is proportional to the standard notion of local phase coherence.
@@
Notation. (1) We will call $\boldsymbol{\theta}$ is an equilibrium of $G$ receives the force $\boldsymbol{\mu}=(\mu_{1},\cdots,\mu_{n})$, if the vetor-labaled graph representation of $\boldsymbol{\theta}$ satisfies (4.1). Note that, in such case, $\mu_{i}=\sum_{j}A_{ij}\cos(\theta_{j}-\theta_{i})$, for all $i\in V$; (2) At state $\boldsymbol{\theta}$, we say that node $i\in V$ is at equilibrium at $\boldsymbol{\theta}$ with respect to $G$ if there exists a scalar $\mu_{i}$ such that (8) holds. Clearly, if $\boldsymbol{\theta}$ is an equilibrium, then every node $i\in V$ is at equilibrium. In particular, we call a node $i$ is positively (or negatively, or zero) forced, if $\mu_{i}>0$ (or <0<0, or =0=0).
4.3. Force and stability
Having established the correspondence between vector alignment at each node and the equilibrium condition, we now turn to the question of stability.
Alignment alone does not guarantee stability. Intuitively, if there exists a node $i$ such that the angle between $\boldsymbol{v}_{i}$ and the resultant vector $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ is zero, a small perturbation will cause the system to return to equilibrium, much like a ball resting at the bottom of a potential valley. In contrast, if the angle is $\pi$, the system is analogous to a ball positioned at the peak of a hill: although it is momentarily at rest, a small perturbation will cause the ball to roll away.
In this section, this intuition is formally confirmed by using the relationship between the negative semi-definiteness of the Jacobian and its diagonal entries, as well as the connection between the stability of equilibrium in gradient dynamical systems and the local geometry of the corresponding energy function.
Figure 6. Illustration of different stability conditions for node $i$. Here, the solid arrow represents $v_{i}$, while the dashed arrow represents $\sum_{j\in N(i)}v_{j}$. The gray region indicates that, in order for node $i$ to attain equilibrium under non-negative forcing, the endpoint of $\sum_{j\in N(i)}v_{j}$ must lie within this sector. (A) Node $i$ is not at equilibrium since (8) does not hold; (B) Node $i$ is at equilibrium but negatively forced, and thus unstable. Therefore, $\boldsymbol{\theta}$ cannot be a second-order critical point. In cases (C) and (D), node $i$ is at equilibrium, and is zero forced and positively forced respectively. Thus, satisfies the necessary condition for $\boldsymbol{\theta}$ to be a second-order critical point.
@@ 4.3. Force and stability
Assume $\boldsymbol{\theta}$ is an equilibrium state of system $G=(V,E)$ or equivalently first order stationary point of (3). In other words, there exist real numbers $\mu_{1},\dots,\mu_{n}$ such that for all $i\in V$,
$\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
If $\boldsymbol{\theta}$ is a second order stationary point of (3), then all $\mu_{i}$ must be non-negative.

### v2

@@ 2. Preliminaries of Threshold Graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines such as scheduling theory, psychology, computer science [threshold_graphs_book]. Threshold graphs have been defined by Chvs and Hammer as follows [ChvatalHammer1973, ChvatalHammer1977].
@@
A graph $G$ is called a threshold graph if it can be constructed from the one-vertex graph by repeatedly applying one of the following two operations:
  1. (i)
    add an isolated vertex, adjacent to no existing vertex;
  2. (ii)
    add a dominating vertex, adjacent to all existing vertices.
The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating.
Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [threshold_graphs_book].
Theorem 2.2.
Let $G=(V,E)$ be a graph. The following statements are equivalent:
  1. (i)
    $G$ is a threshold graph.
  2. (ii)
    $G$ is obtained by iteratively adding isolated or dominating vertices to the one-vertex graph.
  3. (iii)
    $G$ admits a weight function $w:V\to\mathbb{R}$ and a threshold $t\in\mathbb{R}$ such that
    $\{u,v\}\in E\;\Longleftrightarrow\;w(u)+w(v)\geq t,\qquad\forall\,u\neq v.$
  4. (iv)
    $G$ is $\{P_{4},C_{4},2K_{2}\}$-free.
  5. (vi)
    For any distinct $u,v\in V$, $N(u)\subseteq N[v]\quad\text{or}\quad N(v)\subseteq N[u].$
Remark 2.3 (Edge densities of threshold graphs).
The edge density of threshold graphs ranges from $O(1/n)$ to $1$. Indeed, if $(b_{2},\dots,b_{n})$ is the generating sequence, then
$|E|=\sum_{i:b_{i}=1}(i-1),\qquad\mathrm{dens}(G)=\frac{|E|}{\binom{n}{2}}$.
By choosing the positions at which $b_{i}=1$, one obtains connected threshold graphs whose densities range from $2/n$ (stars) to $1$ (the complete graph), spanning many intermediate values. Moreover, threshold graphs are exactly the graphs whose degree sequences saturate the Erdős–Gallai inequalities in the characteristic Ferrers pattern corresponding to nested neighborhoods [threshold_graphs_book].
@@
Figure 3. Forbidden induced subgraphs for threshold graphs.
@@
Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered. To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [Ausiello1977], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs. This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [ChvatalHammer1977].
3. Preliminaries on Kuramoto Model
In this section, we introduce two families of concepts for the Kuramoto model: one arising from dynamical systems and the other from nonconvex optimization. Presenting these notions side by side is important, as many of the local synchronization lemmas developed later rely on interpreting stability to infer the relative positions of phasors at second-order stationary points of the energy landscape.
@@ 3.1. Equilibrium and Local Stability
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [khalil2002nonlinear]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
An equilibrium point is considered stable if small perturbations around it remain small over time [khalil2002nonlinear]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@ 3.2. Stationary Point and Local Optimality
3.3. Correspondence between Stability and Optimality
@@
In this section, we briefly outline the relationship between the equilibrium stability of the dynamical system (2) and the local optimality of the energy function (4). Understanding this correspondence provides valuable intuition. In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order stationary point.
@@
At the first-order level, the two notions coincide. Indeed, a configuration $\boldsymbol{\theta}$ is a first-order stationary point of the energy function (4) if and only if it is an equilibrium of the dynamical system (2) on the graph $G$ with adjacency matrix $A$, since both are characterized by
$\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i})=0\qquad\text{for all }i\in V.$
@@
One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape. It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4). The relation mentioned here is summarized in Figure 4. Such an equivalence was established in the celebrated work [Absil2006StableEquilibrium].
SOSP of 4 Local minima of 4 = stable equilibria of 2
Figure 4. Relation between different concepts.
4. Vector Geometry of Equilibria
In this section, we recast the algebraic first- and second-order stationary conditions in Section 3 into geometric ones. This geometric viewpoint replaces the nonlinear equations (6) and the Hessian constraints by combinatorial relations among the vectors associated with the nodes. They form the starting point of our analysis. Additionally, stronger geometric constraints will be derived in later sections and combined with the results here to rule out all non-synchronous second-order stationary points of the energy function $E_{G}(\boldsymbol{\theta})$ where $G$ is any connected threshold graph.
4.1. Vector-labeled Graph
The basic object underlying this geometric perspective is the vector-labeled graph corresponding to angle configuration of $\boldsymbol{\theta}$. Defining such object allows us reasoning about whether a given state $\boldsymbol{\theta}$ is an equilibrium, and whether an equilibrium is stable, i.e., can we move the angles so as to decrease the energy.
@@ 4.1. Vector-labeled Graph
Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
@@
This is called a phasor, following the standard convention in the Kuramoto model. The object $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$, called vector-labeled graph encoding $\boldsymbol{\theta}$ combinatorially and geometrically.
$\boldsymbol{v}_{i}$$\theta_{i}$
Figure 5. Representation of $\boldsymbol{v}_{i}$ with angle $\theta_{i}$.
@@ 4.2. Alignment and Equilibrium
Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms.
@@
The proof can be found in [monzon2005global], we present here for completeness.
@@
The expression describes a geometric condition of equilibrium: At equilibrium, each phasor aligns with the direction of the sum of its neighbors’ vectors and is scaled to unit length. Although introduced as a normalization factor, the quantity $\mu_{i}$ carries physical meanings. If the neighbors are highly coherent, the sum vector is large, requiring a larger $\mu_{i}$ to normalize $\boldsymbol{v}_{i}$ to unit length; if the directions are dispersed, $\mu_{i}$ is small. In this sense, $\mu_{i}$ measures how strongly the phasors of node $i$’s neighbors align directionally. Indeed, defining the local order parameter $R_{i}:=\frac{1}{\deg(i)}\left|\sum_{j\in N(i)}e^{i\theta_{j}}\right|,$ we have the relation $\mu_{i}=\deg(i)\cdot R_{i}$, showing that $\mu_{i}$ is proportional to the standard notion of local phase coherence.
@@
(A)(B)(C)(D)
Figure 6. Geometric intuition for the first- and second-order stationary conditions. The solid arrow represents $\boldsymbol{v}_{i}$, and the dashed arrow represents $\sum_{j\in N(i)}\boldsymbol{v}_{j}$. For each node $i$, the four panels show all possible configurations satisfying the first-order condition. Among them, cases (A) and (B) violate the second-order condition because the aggregated phasor lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
4.3. Angle and Stability
Given an equilibrium $\boldsymbol{\theta}$, if there exists a node $i\in V$ whose corresponding force satisfies $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point. Indeed, a perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
@@ 4.3. Angle and Stability
If $\boldsymbol{\theta}$ is a second-order stationary point of (4), then either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ or
$\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0.$
Proof of Lemma 4.4.
The $i$-th diagonal entry of $\mathrm{Hess}\,E(\theta)$ can be written as
$\mathrm{Hess}\,E(\boldsymbol{\theta})_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form. If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bhatia2007] [HornJohnson1985]. ∎